OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. (3 points) 49x2 − 28x + 16 9a2 − 30a + 25 25b2 − 45b − 81 16x2 − 24x − 9

Directrix (directrix):

The first and third terms must be perfect squares.

Directrix (directrix):

The middle term must be positive or negative twice the product of the square roots of the first and third terms. First, middle, and last terms refer to the polynomial in descending order of variable.

OpenStudy (anonymous):

thank you i wasnt sure how he came up with that answer

Directrix (directrix):

Look for the first and third terms as perfect squares first. Then, see about the middle term. If the answer is not B, what do you think it is.

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

no c?

Directrix (directrix):

Let me check B again. Wait a second.

Directrix (directrix):

9a^2 − 30a + 25 --------------- 1. 9a^2 = (3a)^2 2. 25 = 5^2 3. Does -30a = positive/negative 2* (3a)*5?

OpenStudy (anonymous):

?

Directrix (directrix):

3. Does -30a = positive/negative 2* (3a)*5? Yes. So my initial response is incorrect. Let me check the other options.

Directrix (directrix):

9a^2 − 30a + 25 = (3a - 5) ^2 B is correct.

OpenStudy (anonymous):

okay thank you

Directrix (directrix):

Glad to help. Sorry I messed up at the start.